# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v2

@@ 1 Introduction
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL23+] that when $s<\sqrt{\alpha}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
    Can we provide the evidences that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there are evidences suggesting that detection is impossible (by efficient algorithms)?
@@
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold, there are evidences suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold, we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
@@ 1.1 Key challenges and innovations
However, there are certain obstacles when implementing the above ideas, as we shall discuss below. In Step (2) we need to “transfer” the behavior of a statistics $g_{i}$ under $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ to its behavior under $\mathbb{Q}$. If the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by $1+o(1)$, then the standard low-degree conjecture implies that we cannot distinguish between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ better than random by efficient algorithms. Consequently, we expect the behavior of $g_{i}$ to be “almost identical” under both $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ (as $g_{i}$ can be efficiently computed). However, in our case the low-degree advantage is just bounded by a large (but fixed) constant. This weaker condition means that the standard low-degree conjecture only rules out the possibility to distinguish these two measures efficiently with vanishing errors, leaving room for non-negligible distinctions. Similar issues also arise in Step (4), in which simply showing the low-degree advantage between $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ and the low-degree advantage between $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ is bounded is not strong enough for our goal.
To address these issues, one of the main conceptual contributions in our work is to give a more refined characterization of the limitations of efficient algorithms when the low-degree advantage is $O(1)$. Specifically, we show that (assuming low-degree conjecture) bounded low-degree advantage does not only excludes all algorithms that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$, but also suggest a certain kind of algorithmic contiguity. To be more precise, if the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}$ is bounded, then there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}=1)=\Omega(1)$. This framework allows us to transfer the behavior of the behavior of a efficiently computable statistics $g$ under different probability measures more easily. For example, if we know that $\mathbb{P}(g\geq c)\geq\Omega(1)$ it immediately holds that $\mathbb{Q}(g\geq c)\geq\Omega(1)$.
@@ 1.2 Comparison to concurrent work
However, the reduction technique employed in this paper differs significantly from that of [DHSS25+]. The authors of [DHSS25+] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward as we leverage the universal framework described in Subsection 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25+], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
@@ 2 Low-degree framework and algorithmic contiguity
The low degree conjecture, proposed in [Hop18], can be summarized as follows.
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements are equivalent:
@@
    $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$ (respectively, $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=1+o(1)$) as $n\to\infty$.
@@
    There is no algorithm with running time $n^{\widetilde{O}(D)}$ that achieves strong detection (respectively, weak detection) between $\mathbb{P}$ and $\mathbb{Q}$.
@@
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Subsection 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity.
@@
We say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ is a $\mathbb{Q}$-based one-sided test, if
@@
We say that $\mathbb{P}$ is degree-$D$ algorithmic contiguous with respect to $\mathbb{Q}$, denoted as $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$, if no $\mathbb{Q}$-based one-sided testing algorithm runs in time $n^{\widetilde{O}(D)}$. We say that $\mathbb{Q}$ and $\mathbb{P}$ are degree-$D$ algorithmic mutually contiguous, denoted as $\mathbb{Q}\bowtie_{\leq D}\mathbb{P}$, if both $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$ and $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$ hold.
@@
Assuming the Conjecture 2.2, for the high-dimensional hypothesis testing problem between $\mathbb{P}$ and $\mathbb{Q}$, if $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ and $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then we have $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$.
The proof of Theorem 2.4 is incorporated in Subsection B.1 of the appendix. In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Subsection 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture, which posits that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements are equivalent:
@@
    $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ (respectively, $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$) as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$.
@@
    There is no algorithm with running time $n^{\widetilde{O}(D)}$ that achieves strong detection (respectively, weak detection) between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    For all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{\widetilde{O}(D)}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$ (respectively, $\mathsf{Adv}(f)=1+o(1)$), where
@@ 3 Partial recovery in correlated random graphs
    For the correlated Erdős-Rényi graphs $\mathcal{G}(n,q,\rho)$ where $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There is no algorithm with running time $n^{\widetilde{O}(D)}$ that achieves weak partial matching, provided that
@@
    For the correlated SBMs $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ where $\lambda=O(1)$ and $\epsilon^{2}\lambda<1-\delta,s<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There is no algorithm with running time $n^{\widetilde{O}(D)}$ that achieves strong partial matching, provided that
    ${}D=O(\log n)\,.$ (3.2)
@@
    If $(G,A,B)\sim\mathcal{G}(n,q,\rho)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $q,\rho,D$ satisfy the assumptions in Item (1) of Theorem 3.2. Then for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}]=o(1)$ for all $1\leq i\leq n$.
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
@@
The proof of Lemma 3.5 is quite technical and thus we postpone it to Subsection B.2 of the appendix. Now, based on Lemma 3.5, we establish the following result, which basically suggests that it is impossible to obtain a “good approximation” of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$ in some sense.
@@
Assuming Conjecture 2.5, we have for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$, it holds that
@@
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ such that
@@
Thus, using the equivalence of Items (2) and (3) in Conjecture 2.5 on the testing problem between $\mathbb{P}$ and $\mathbb{Q}$, we see that there exists an algorithm with running time $n^{\widetilde{O}(D)}$ that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$. However, it is impossible assuming Conjecture 2.5 as [DDL23+] showed that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$ where $\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P})=o(1)$. This implies that (3.6) is impossible and thus completes our proof. ∎
@@
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. Using Items (1) and (2) in Definition 3.1, we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
@@
Similarly as in Subsection 3.1, our first step is to show the following lemma.
@@
The proof of Lemma 3.9 is incorporated in Subsection B.3 of the appendix. Based on Lemma 3.9, we can deduce our main result just as how we deduce Theorem 3.2 from Lemma 3.5. The only difference is that we will replace all $\overline{\mathbb{P}}$ with $\widetilde{\mathbb{P}}$ and replace all $\mathbb{P}$ with $\mathbb{P}(\cdot\mid\mathcal{E})$ so we omit further details here.

### v3

@@ 1 Introduction
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL23+] that when $\rho<\sqrt{\alpha}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
    Can we provide the evidence that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there are evidence suggesting that detection is impossible (by efficient algorithms)?
@@
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold (i.e., when $\epsilon^{2}\lambda s<1$), there is evidence suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold (i.e, when $\epsilon^{2}\lambda s>1$), we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
@@ 1.1 Key challenges and innovations
However, there are certain obstacles when implementing the above ideas, as we shall discuss below. In Step (2) we need to “transfer” the behavior of a statistic $g_{i}$ under $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ to its behavior under $\mathbb{Q}$. If the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by $1+o(1)$, then the standard low-degree conjecture implies that we cannot distinguish between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ better than random by efficient algorithms. Consequently, we expect the behavior of $g_{i}$ to be “almost identical” under both $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ (as $g_{i}$ can be efficiently computed). However, in our case the low-degree advantage is just bounded by a large (but fixed) constant. This weaker condition means that the standard low-degree conjecture only rules out the possibility to distinguish these two measures efficiently with vanishing errors, leaving room for non-negligible distinctions. Similar issues also arise in Step (4), in which simply showing the low-degree advantage between $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ and the low-degree advantage between $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ is bounded is not strong enough for our goal.
To address these issues, one of the main conceptual contributions in our work is to give a more refined characterization of the limitations of efficient algorithms when the low-degree advantage is $O(1)$. Specifically, we show that (assuming low-degree conjecture) bounded low-degree advantage does not only exclude all algorithms that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$, but also suggest a certain kind of algorithmic contiguity. To be more precise, if the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}$ is bounded, then there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}=1)=\Omega(1)$. This framework allows us to transfer the behavior of a efficiently computable statistics $g$ under different probability measures more easily. For example, if we know that $\mathbb{P}(g\geq c)\geq\Omega(1)$ it immediately holds that $\mathbb{Q}(g\geq c)\geq\Omega(1)$.
@@ 1.2 Comparison to concurrent work
However, the reduction technique employed in this paper differs significantly from that of [DHSS25+]. The authors of [DHSS25+] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward as we leverage the universal framework described in Section 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25+], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
@@ 2 Low-degree framework and algorithmic contiguity
Note that if we denote the likelihood ratio $L(\mathsf{Y})=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}(\mathsf{Y})$, it is well-known (see, e.g., [Hop18]) the right hand side equals the $L_{2}$-norm of the projection of $L(\mathsf{Y})$ into the subspace spanned by all polynomials of degree bounded by $D$ (the norm is induced by natural inner product under $\mathbb{Q}$) and thus we characterize it with $\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}$. The low degree conjecture, proposed in [Hop18], can be summarized as follows.
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=1+o(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Section 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity.
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, we say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ is a $\mathbb{Q}$-based one-sided test, if
@@
We say that $\mathbb{P}$ is time-$D$ algorithmic contiguous with respect to $\mathbb{Q}$, denoted as $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$, if no $\mathbb{Q}$-based one-sided testing algorithm runs in time $n^{D}$. We say that $\mathbb{Q}$ and $\mathbb{P}$ are degree-$D$ algorithmic mutually contiguous, denoted as $\mathbb{Q}\bowtie_{\leq D}\mathbb{P}$, if both $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$ and $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$ hold.
@@
Assuming the Conjecture 2.2, for the high-dimensional hypothesis testing problem between $\mathbb{P}$ and $\mathbb{Q}$, if $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ and $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then we have $\mathbb{P}\lhd_{\leq D/(\log n)^{C}}\mathbb{Q}$ for some constant $C$.
The proof of Theorem 2.4 is incorporated in Section B.1 of the appendix. In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Section 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture, which posits that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    If there is no algorithm with running time $n^{D}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$, where
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
  4. (4)
    If there is no algorithm with running time $n^{D}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=1+o(1)$).
@@ 3 Partial recovery in correlated random graphs
    For the correlated Erdős-Rényi graphs $\mathcal{G}(n,q,\rho)$ where $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak partial matching, provided that
@@
    For the correlated SBMs $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ where $\lambda=O(1)$ and $\epsilon^{2}\lambda<1-\delta,s<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong partial matching, provided that
    ${}D=n^{o(1)}\,.$ (3.2)
@@
    If $(G,A,B)\sim\mathcal{G}(n,q,\rho)$ and let $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $q,\rho,D$ satisfy the assumptions in Item (1) of Theorem 3.2. Then there exists a constant $C$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}]=o(1)$ for all $1\leq i\leq n$.
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ and $D$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists a constant $C$ and an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$, and for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
@@
The proof of Lemma 3.5 is quite technical and thus we postpone it to Section B.2 of the appendix. Now, based on Lemma 3.5, we establish the following result, which basically suggests that it is impossible to obtain a “good approximation” of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$ in some sense.
@@
Assuming Conjecture 2.5, there exists constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, it holds that
@@
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that
@@
Thus, using Item (2) in Conjecture 2.5 on the testing problem between $\mathbb{P}$ and $\mathbb{Q}$, we see that there exists an algorithm with running time $n^{D/(\log n)^{C}}$ that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$. However, it is impossible assuming Item (1) of Conjecture 2.5 as [DDL23+] showed that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$ where $\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P})=o(1)$. This implies that (3.6) is impossible and thus completes our proof. ∎
@@
We choose $C$ as in the statement of Lemma 3.6. Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. Using Items (1) and (2) in Definition 3.1, we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
@@
Similarly as in Section 3.1, our first step is to show the following lemma.
@@
The proof of Lemma 3.9 is incorporated in Section B.3 of the appendix. Based on Lemma 3.9, we can deduce our main result just as how we deduce Theorem 3.2 from Lemma 3.5. The only difference is that we will replace all $\overline{\mathbb{P}}$ with $\widetilde{\mathbb{P}}$ and replace all $\mathbb{P}$ with $\mathbb{P}(\cdot\mid\mathcal{E})$ so we omit further details here.